# Toward A Reinforcement-Learning-Based System for Adjusting Medication to Minimize Speech Disfluency

### 12 Dec 2023 v1

@@ Results
where $S$ is current speech fluency, $D$, $A$, $I$ are the patient’s current depression, anxiety, and insomnia scores respectively, labelled on a 5 point scale, where 1 represents a patient who does not suffer from the given effect, and 5 represents a patient who suffers from it in it’s most malignant form. This data can be collected through regular patient check-ups and be formed as a scale for each individual. $R$ is the patient’s regular speech fluency score, a value that may be different for every person, and was computed as a weighted sum utilizing factors such as age. Our implementation of the LinUCB algorithm observes the current state of the patient, then for each medication that is part of the environment, it estimates the reward using a linear approximation. An upper confidence bound is calculated for the estimated reward in which the medication with the highest upper confidence bound is chosen according to the equation: $a_{t}=argmax_{a\epsilon A}(x_{t}(\alpha)^{T}\theta_{a}+\alpha\sqrt{x_{t}(a)^{T}A_{a}^{-1}x_{t}(a)})$ where $x_{t}$(a) is the feature vector for action $a$ at time $t$, $\theta_{a}$ is the parameter vector for action $a$ which we want to estimate, and $\alpha$ is the hyperparameter controlling how exploitative/exploratory the agent wants to be in making decisions. The higher the value of $\alpha$, the more the agent will venture out and choose new medications, the lower the value of $\alpha$ the more patient will choose medications it has current knowledge of. In our implementation of the LinUCB algorithm, we chose the value of $\alpha=10.0$ to get our results, as it demonstrated an acceptable bound for the algorithms exploratory tendency and provided a representative time frame for the onset of speech disfluency to decrease. For training our algorithm, we set up a training loop simulating a single day where the agent uses a dynamic step driver responsible for executing the agent’s policy in the environment for a specified number of steps. A step in our environment consists of administering the selected medication to the patient and subsequently updating the states of the patient based on the effects of the medication. After each step it collects data for the agent which is stored in a replay buffer which allows for efficient learning by sampling from these experiences.
Results
In Figure 3, we see the 3 most common outcomes of the medication selection algorithm. In the top graph, we see a steady increase in the speech fluency of a patient. The model converges, and we see that the model has success in selecting accurate medications. The graph in the middle demonstrates an initial increase and then a plateau in the medication selection. This is indicative of the model struggling to find the most suitable medication for the patient, resulting in a prolonged period of uncertainty and potentially ineffective treatment. Finally, in the bottom graph, we observe a sharp decline in speech fluency, suggesting that the medication selection algorithm failed to provide effective treatment for the patient, leading to worsening symptoms. These three distinct outcomes illustrate the variability in the medication selection process and emphasize the importance of continuous monitoring and adjustment of the algorithm to improve patient outcomes.
Figure 3: Speech Fluency over time
We ran a limited set of simulations where the medications had plausible properties, and the precision of the measurement of the speech disfluency was to within 10% (note that we obtained results indicating that this is possible). Our preliminary results indicate that if speech disfluency can be measured to within 10% of the true score and the medications have plausible properties (similar to the ones seen in Medication Literature Review), reinforcement learning is possible to pessimize speech disfluency.
Medication Literature Review
@@
A systematic literature review was conducted to determine common medications used to treat major depressive disorder (“depression”) and other mental illnesses that affect speech fluency. Table LABEL:t1 indicates 23 medications whose onset time and response rates were used to inform the reinforcement learning simulation.
Ethical Considerations
@@
In this paper, we outline and evaluate a proposal to adjust medications automatically in order to improve the speech fluency of a simulated patient. Although administering medications automatically is sometimes done (e.g., with Insulin pumps), this can only be done after thorough clinical trials and with informed consent from the patient. Patients must be thoroughly informed about the nature of the automated system, its potential risks and benefits, and their rights in the decision-making progress. This extends beyond initial consent, and includes ongoing consent as the system adapts medication regiments. We believe that if safe dosage ranges can be determined, automatically adjusting medications can in principle be done. Potential upsides include better-optimized medication regimes that are found faster (or found at all). The system’s reliance on sensitive patient data raises significant privacy and security concerns. Strict adherence to data protection laws and ethical guidelines is mandatory. Finally, the long-term implications of integrating such a system into healthcare practices must be considered. This includes ongoing monitoring of the system’s impact on patient health outcomes, adapting to new medical insights, and updating the system accordingly.
Limitations
@@
This paper studies the feasibility of using automated measures of disfluency together with a reinforcement learning system to adjust medications. Our model is informed by the literature, but it is still a “toy” model. However, our results indicate that it is possible to measure disfluency to a certain precision (for some definition of “disfluency”), and that this precision can be sufficient to serve as input to an RL system.
Conclusions and Future Work
@@
We propose and evaluate a new idea: people with speech disfluency can have their disfluency measured automatically, and a Reinforcement Learning algorithm can find an optimal medication regime for them through exploration/exploitation. We have shown that the components of this system are possible to build. In our preliminary experiments, we have shown that with an accurate enough disfluency detection system and with medications with plausible properties, a good medication regime can be found by a Reinforcement Learning system. Future work includes improving our subsystems for performance, ensuring that our simulations preserve patient safety (e.g. by only trying safe dosages), and running more extensive simulations to show under what conditions RL for optimizing medications is feasible.

### 06 Feb 2024 v4

@@ 5.6 Medication Selection Algorithm
where $S$ is current speech fluency, $D$, $A$, $I$ are the patient’s current depression, anxiety, and insomnia scores respectively, labelled on a 5 point scale, where 1 represents no symptoms and 5 represents the most severe symptoms. $R$ is the patient’s baseline fluency.
The implementation of the LinUCB algorithm observes the current state of the patient, then for each medication that is part of the environment, it estimates the reward using a linear approximation. An upper confidence bound is calculated for the estimated reward in which the medication with the highest upper confidence bound is chosen according to the equation:
$a_{t}=\arg\max_{a\epsilon A}\left(x_{t}(\alpha)^{T}\theta_{a}+\alpha\sqrt{x_{t}(a)^{T}A_{a}^{-1}x_{t}(a)}\right)$
where $x_{t}(a)$ is the feature vector for action $a$ at time $t$, $\theta_{a}$ is the parameter vector for action $a$ which we want to estimate, and $\alpha$ is the hyperparameter controlling the exploration/exploitation trade-off [20]. In our implementation of the LinUCB algorithm, we chose the value of $\alpha=10.0$.
5.7 Results
In our simulations, we run the RL algorithm and keep track of disfluency over time. We inject noise into the algorithm’s simulated measurements of disfluency to simulate the fact that our disfluency detection system does not measure disfluency perfectly. In the experiments reported here, we inject a minimal amount of noise, corresponding to the high precision with which we can measure disfluency.
We define the success of a trial as an improvement of over $0.5\sigma$ in speech fluency. We define failure as a deterioration of over $0.5\sigma$ in speech fluency. Here, $\sigma\approx 0.1$ is the standard deviation of fluency in the dataset.
Figures 5.7 and 5.7 show examples of a successful and an unsuccessful run, respectively. Across 500 patient simulation runs, we found a high potential for reinforcement learning to correctly apply medication effects to reduce speech disfluency, with 52% of runs showing success and 9% of runs demonstrating failure.
figureConvergence to higher speech fluency.
figureLack of convergence to higher speech fluency.
The average fluency across these runs was 0.66/1.00 with a standard deviation of 0.1. The success rate of the simulations was 52%, with a failure rate of approximately 16%. Success and failure are defined as runs terminating greater than $0.5\sigma$ and lower than .$-.5\sigma- .$ from the initial fluency level, respectively.
These results support the possibility of the use of reinforcement learning to improve speech fluency under the studied conditions. Our preliminary results indicate that if speech disfluency can be measured to within 10% of the true score and the medications have plausible properties (similar to the ones seen in Section 6), then reinforcement learning is a possible method to dose medications that pessimize speech disfluency. However, the variability in outcomes and the presence of failed simulations that prompted theoretical patient deterioration indicate that further research is needed in improving the model’s accuracy and understanding factors contributing to failures, that will be important for applying these findings in a clinical setting.
6 Medication Literature Review
@@
A systematic literature review was conducted to determine common medications used to treat major depressive disorder (“depression") and other mental illnesses that affect speech fluency. Table 1 indicates 23 medications whose onset time and response rates were used to inform the reinforcement learning simulation.
7 Ethical Considerations
@@
In this paper, we outline and evaluate a proposal to adjust medications automatically in order to improve the speech fluency of a simulated patient. Although administering medications automatically is sometimes done (e.g., with Insulin pumps), this can only be done after thorough clinical trials and with informed consent from the patient. Patients must be thoroughly informed about the nature of the automated system, its potential risks and benefits, and their rights in the decision-making progress. This extends beyond initial consent, and includes ongoing consent as the system adapts medication regimens. We believe that if safe dosage ranges can be determined, automatically adjusting medications can in principle be done. Potential upsides include better-optimized medication regimes that are found faster (or found at all). The system’s reliance on sensitive patient data raises significant privacy and security concerns. Finally, the long-term implications of integrating such a system into healthcare practices must be considered. This includes ongoing monitoring of the system’s impact on patient health outcomes, adapting to new medical insights, and updating the system accordingly.
8 Limitations
@@
This paper studies the feasibility of using automated measures of disfluency together with a reinforcement learning system to adjust medications. Our model is informed by the literature, but it is still a “toy" model. However, our results indicate that it is possible to measure disfluency to a certain precision (for some definition of “disfluency"), and that this precision can be sufficient to serve as input to an RL system.
9 Conclusions and Future Work
@@
We propose and evaluate a new idea: people with speech disfluency can have their disfluency measured automatically, and a reinforcement learning algorithm can find an optimal medication regime for them through exploration/exploitation. We have shown that the components of this system are possible to build.
In our preliminary experiments, we have shown that with an accurate enough disfluency detection system and with medications with plausible properties, a good medication regime can be found by a reinforcement learning system.
Future work includes improving our subsystems for performance, ensuring that our simulations preserve patient safety (e.g. by only trying safe dosages), and running more extensive simulations to show under what conditions RL for optimizing medications is feasible.